OpenStudy (anonymous):

how do you get the length of a copper wire if the diameter is .10cm the mass is 3.4449 g and the density is 8.96g/cm^3

OpenStudy (anonymous):

find the cross-sectional area. volume is length*cross-sectional area volume*density = mass volume = mass/density length = mass/(density*cross-sectional area)

OpenStudy (ghazi):

cross sectional area= pi*(10/2)^2

OpenStudy (ghazi):

volume = pi* r^2*length

OpenStudy (ghazi):

volume= mass/ density

OpenStudy (ghazi):

\[volume=\frac{ 3.4449 }{ 8.96 }=0.38449 cm^3\] now \[0.38449= \pi* 5^2*L\] and \[L=\frac{ 0.38449}{ 3.14*25 }=0.0048 \]

OpenStudy (phi):

@crystal9193 @ghazi has the correct approach, but used the incorrect diameter of 10 cm rather than 0.10 cm use r= 0.05 cm in pi*r^2 to find the cross-sectional area of the wire.